OpenStudy (anonymous):

(16x^2 +4x +20) / 4x Help, please.

OpenStudy (anonymous):

divide term by term

OpenStudy (anonymous):

\[\frac{16x^2}{4x}=4x\] for example

OpenStudy (anonymous):

So the complete answer would be 4x + 5/x ?

OpenStudy (anonymous):

(16x^2 +4x +20) / 4x \[\frac{16x^{2}+4x+20}{4x}=\frac{16x^{2}}{4x}+\frac{4x}{4x}+\frac{20}{4x}\]

OpenStudy (anonymous):

Yes. Correct...

OpenStudy (anonymous):

so just simplify.to what you had.

OpenStudy (anonymous):

4x = 5/x ?

OpenStudy (anonymous):

just multiply both sides by x divide both sides by four and then take the square root of both sides. remember +-